OpenStudy (anonymous):

Can someone show me how to solve this? I'm confused because I don't know how to solve it when there is a square root on the top and bottom 15√6 ------ 3√5

OpenStudy (anonymous):

I'm supposed to rationalize and simplify

jhonyy9 (jhonyy9):

so try it what you have supposed

jhonyy9 (jhonyy9):

courage

OpenStudy (anonymous):

I don't know how to rationalize and simplify when there is a square root thing on the top and the bottom @jhonyy9

OpenStudy (johnweldon1993):

Try multiplying both the top and bottom by squareroot5

OpenStudy (anonymous):

okay, so I got 15√30 is that right? ------ 8

OpenStudy (johnweldon1993):

Not quite..what happened on the bottom there? \[3\sqrt{5} \times \sqrt{5} = 3 * 5= 15\]

jhonyy9 (jhonyy9):

how you have got this 8 ?

OpenStudy (anonymous):

but aren't you supposed to multiply the √5 by the other √5 and the that would be √25 witch would equal 5 and 3+5=8

OpenStudy (johnweldon1993):

All of this is multiplication,where did that + sign come in ?

OpenStudy (johnweldon1993):

\[\large \frac{15\sqrt{6}}{3\sqrt{5}} \times \frac{\sqrt{5}}{\sqrt{5}} = \frac{15\sqrt{6}\sqrt{5}}{3\sqrt{5}\sqrt{5}} = \frac{15 \sqrt{30}}{3 \times 5} = \frac{15\sqrt{30}}{15}\] Right?

OpenStudy (anonymous):

I'm not seeing how 3√5 (√5) is 15.... shouldn't the fives be multiplied together? 3√5(√5) -> 3√25 -> 3√5(5) -> 3+5=8

OpenStudy (johnweldon1993):

Alright, so from \[\large 3\sqrt{5}\sqrt{5} = 3\sqrt{5\times 5} = 3\sqrt{25}\] And yes...the square root of 5 IS 5...but why are you putting a '+' sign? right now it is 3 TIMES the square root of 25...so when that changes to 5 it will be 3 TIMES 5 ...there is no '+' sign added when you simplify a square root \[\large 3\sqrt{25} = 3 \times 5 = 15\]

OpenStudy (anonymous):

ohhhh I see it now! so now that I have 15√30 can I simplify it more? ----- 15

OpenStudy (johnweldon1993):

Yes absolutely!

OpenStudy (anonymous):

so then I would divide it by 15?

OpenStudy (johnweldon1993):

correct! don't turn it into a decimal though that would be messy..I would just keep it as the radical

OpenStudy (anonymous):

okay then, tell me if i'm doing this right... 15√30 ------- = √30? 15

OpenStudy (johnweldon1993):

That's perfect! \[\frac{\cancel{15}\sqrt{30}}{\cancel{15}} = \sqrt{30}\]

OpenStudy (anonymous):

Thanks so much for the help!

OpenStudy (johnweldon1993):

No problem!